# 2024_subroc.bib

### v1

@@ 3.1.3 Generalization Awareness
It is a common issue in subgroup discovery that multiple subgroups in the result set have a high pairwise overlap in terms of their cover [lemmerich_difference-based_2013]. To reduce the intersections of subgroup covers in result sets, there is an adaptation to scoring functions available from literature, called generalization awareness. For this, a variant [lemmerich_difference-based_2013] of previous approaches[batal_concise_2010, lemmerich_local_2011, grosskreutz_subgroup_2010] extends a scoring function $\varphi^{\prime}$ by reducing the score $\varphi^{\prime}(p)$ of a subgroup with pattern $p$ by subtracting the maximum score across all its generalizations $h$ (making a specialization viable only if its score is substantially higher):
@@ 3.2 Optimistic Estimates
For runtime optimization, we derive tight optimistic estimates for the scoring functions $\varphi^{rasl}$, $\varphi^{rROCAUC}$ and $\varphi^{rPRAUC}$ as well as, separately, for the cover size and class balance factors. Proofs are given in Appendix A.
@@ 3.2.1 Average Ranking Loss
To date, no (tight) optimistic estimate was derived for the ARL based scoring function $\varphi^{rasl}$ [duivesteijn_understanding_2014]. The $ARL(p)$ is defined as the average over the ranking losses $PEN(c,p)$ of all covered positive instances $c\in\mathcal{P}_{p}$. The penalty $PEN(c,p)$ for each positive instance $c$ counts the negative instances ranked higher than $c$. Thus, removing positive instances with non-maximal $PEN(c,p)$ does not alter the $PEN$ of the remaining instances and increases the average by eliminating small values. The $ARL(p)$ is therefore maximized when only a single positive $c^{\prime}$ instance with maximal $PEN(c^{\prime},p)$ is left.
@@ 3.2.2 Area Under the Receiver-Operating Curve
Next, we introduce a lower bound $b$ for ROC AUC that we will use to derive an optimistic estimate of $\varphi^{rROCAUC}$. We distinguish three cases: (a) If perfect separation of ground truth labels is possible with a threshold on the soft predictions, the same holds for every subset. Thus, the ROC AUC is always 1. (b) If the input additionally contains some instances that share the same prediction value (ties) but have different labels (positive and negative), then the ROC AUC on only these tied instances equals $\frac{1}{2}$. Any subset of the given instances containing both classes and at least one non-tied instance allows for the correct separation of some positives and negatives by at least one threshold. Thus, the ROC AUC is greater than $\frac{1}{2}$ for such subsets. (c) The input contains at least one incorrectly ranked pair of a positive and a negative instance. The subset containing exactly one such pair has an ROC AUC of 00, which is the lowest possible ROC AUC.
@@ 3.2.3 Area Under the Precision-Recall Curve
Lemma 4 introduces a lower bound for PR AUC which is used in Theorem 5 to derive a tight optimistic estimate for $\varphi^{rPRAUC}$. For Lemma 4, a subset $I_{worst}$ of the original set of instances $I$ is constructed to maximize classification errors while minimizing the positive class ratio. This makes it impossible for a threshold on any other subset to define a point in PR space below the PR curve of $I_{worst}$ (with some exceptions that have no impact on the AUC of the resulting PR curve). Proofs and an illustration are included in Section A.3.
@@
3.2.4 Cover Size and Class Balance Weighted Scoring Functions
The following Lemma 6 gives a tight upper bound for the weighting used in cover size and class balance weighted scoring functions when $\alpha\leq\beta$. This result translates to optimistic estimates for $\varphi_{\alpha,\beta}^{rasl}(\cdot)$, $\varphi_{\alpha,\beta}^{rROCAUC}(\cdot)$ and $\varphi_{\alpha,\beta}^{rPRAUC}(\cdot)$. It therefore enables the use of the optimistic estimates presented in theorems 1, 3 and 5 when cover size and class balance weighting is enabled.
@@ 3.3 Significance Filtering
Subgroup discovery inspects an exponential search space of patterns and consequently requires correction for multiple hypothesis comparison. To reduce the number of compared subgroups we use a holdout procedure [webb_discovering_2007] splitting analyzed datasets into two parts, the search dataset to search subgroups and the validation dataset to test statistical significance. The tested statistic is the raw score of a subgroup (written $\varphi_{0,0}$) excluding cover size and class imbalance. The null hypothesis is that $\varphi_{0,0}(p)$ for a pattern $p$ is drawn from the sample distribution of $\varphi_{0,0}$ on random instances with the same size and class balance as the cover $sg(p)$ of $p$. This allows a test on performance differences, excluding the influence of cover size and class balance on the used performance measure.
We compute the empirical p-value of a subgroup $p$ by a randomization test (cf. [gionis_assessing_2007]), where $1000$ random subsets of the validation dataset are drawn without replacement and with fixed size and class balance. $\varphi_{0,0}$ is evaluated on each of the random subsets to construct an empirical distribution of exceptionality. The empirical p-value is the fraction of random subsets with interestingness $\varphi_{0,0}$ greater than or equal to $\varphi_{0,0}(p)$ divided by the total number of random subsets.
Multiple testing correction is done using Benjamini-Yekutieli (BY) [benjamini_control_2001] (controlling the false discovery rate (FDR)). BY has the necessary properties to be applied under any dependency between tested hypotheses. The correction approach can be customized by the user, e.g., with Bonferroni controlling familywise error rate (FWER). The default p-value threshold we use is $p\leq 0.05$.
@@ 4 Experiments
Datasets are from OpenML [vanschoren_openml_2013] and the UCI Machine Learning Repository [kelly_uci_nodate] (statistics in Supp. Table 3). Each was partitioned randomly into four equal parts for training, subgroup discovery, significance testing, and result evaluation.
4.1 Quality of SubROC Results Across Multiple Datasets
To illustrate the quality of subgroups discovered with SubROC, we applied xgboost on various datasets and searched for exceptionally underperforming subgroups with regard to the performance measures ROC AUC, PR AUC, and ARL. Table 1 compares two settings: a basic SubROC setting (b) and the full SubROC framework (f). For b, the search solely uses the scoring functions $\varphi^{rasl}$, $\varphi^{rROCAUC}$, and $\varphi^{rPRAUC}$. The top-5 subgroups are returned and additionally tested for significance. The full framework (f) is configured with size and class imbalance weighting $\alpha=\beta=1$, generalization awareness, and statistical significance filtering. Here, the search returns the top-100 subgroups, which were filtered for significance, and finally, the top-5 subgroups are returned. In both cases, we used a minimum subgroup size of $\geq 20$, and a maximum pattern length of 4. Statistics were calculated for top-5 results. Note that exceptionality should not be compared between different performance measures because of their differing scales and responses to errors.
For b, cover sizes tend to be close to the minimum cover size limit, and class ratios tend to be close to 0 or 1. Despite their high difference in performance compared to the overall dataset (exceptionality), such subgroups do not necessarily point towards generalizable structural issues of the classifier but may be caused by artifacts (cf. Section 4.4). This is also apparent in the low number of subgroups passing statistical significance in the top-5 results of b. Similar patterns are observable in the top-100 results (see Supp. Table 5). While yielding less prominent exceptionality, we observe that the SubROC framework leads to larger and more balanced results. Additionally, substantially more subgroups are returned that pass the significance filtering step. This is apparent in the top-5 as well as the top-100 result (cf. Supp. Table 5).
We also inspect the pairwise Intersection over Union (IoU) of the discovered top-5 subgroups. A higher IoU refers to more overlapping and, thus, redundant subgroups. Values generally remain in a similar range to the baseline, but we observe occasional significant improvements by SubROC. This indicates that our framework does not come at a cost of and in some cases improves result diversity.
The statistical filtering step of SubROC saves resources for practitioners by discarding a considerable amount of initially found, but unreliable subgroups. This effect is particularly prominent for small datasets with less than $2,500$ search instances, where the significance filtering discarded almost all subgroups. It is less prominent for ARL based scoring functions, which generally produce smaller result sets with a greater share of significant subgroups. Additional results can be found in Section B.2. For concrete result sets from a case study on the Adult dataset see Section B.3.
Table 1: Overview of top-5 result set properties from a basic SubROC setting (b) and the full SubROC framework (f) for a range of datasets and performance measures. SubROC managed to find diverse and exceptional subgroups, which are often much larger and balanced than the baseline. (Continued)
@@ 4.1 Quality of SubROC Results Across Multiple Datasets
Top-5/ Significant/ Original
Dataset Metric     b f     b f
Census n=199523 NCR=0.06 ROC     25 6753    1/5 5/99/100
PR     23 864    0/5 5/99/100
ARL     8000 11945    5/5 5/27/27
Adult n=48842 NCR=0.24 ROC     27 6492    0/5 5/77/100
PR     27 2803    2/5 5/70/100
ARL     4614 6930    5/5 5/5/5
Bank n=45211 NCR=0.88 ROC     22 311    0/5 5/22/100
PR     157 72    0/5 2/2/100
ARL     - -    0/0 0/0/0
Credit n=30000 NCR=0.78 ROC     24 1506    0/5 5/38/100
PR     75 1895    0/5 5/17/100
ARL     - -    0/0 0/0/0
Approval n=690 NCR=0.55 ROC     29 -    0/5 0/0/41
PR     41 80    2/5 1/1/25
ARL     60 87    3/5 3/3/3
Table 1: Continued from previous columns. (Rows continued in Supp. Table 4, containing results for more datasets)
Mean Exceptionality
@@
Dataset Metric     b f     b f     b f
Census n=199523 NCR=0.06 ROC     0.95 0.05    0.04 0.16    0.10 0.31
PR     0.92 0.14    0.93 0.45    0.37 0.45
ARL     47 31    0.15 0.11    0.99 0.50
Adult n=48842 NCR=0.24 ROC     0.92 0.06    0.04 0.37    0.48 0.61
PR     0.94 0.08    0.96 0.46    0.37 0.19
ARL     148 55    0.42 0.34    0.85 0.61
Bank n=45211 NCR=0.88 ROC     0.83 0.16    0.95 0.48    0.10 0.12
PR     0.68 0.10    0.99 0.48    0.80 0.51
ARL     - -    - -    - -
Credit n=30000 NCR=0.78 ROC     0.75 0.12    0.96 0.61    0.01 0.09
PR     0.51 0.29    0.99 0.85    0.18 0.34
ARL     - -    - -    - -
Approval n=690 NCR=0.55 ROC     0.64 -    0.96 -    0.69 -
PR     0.89 0.85    0.98 0.95    0.75 -
ARL     26 10    0.96 0.79    0.68 0.69
4.2 Optimistic Estimates Substantially Speed Up Runtimes
To investigate the efficiency improvements enabled by SubROC, we measure runtimes of subgroup discovery with and without pruning based on the introduced optimistic estimates 10 times for each dataset (3 times for Census due to long runtimes). The searches run on predictions from a Gaussian Naive Bayes classifier, using the best-first-search enumeration strategy, with a maximum pattern length of 4. Different size and class imbalance weightings ($\alpha,\beta$) are compared. Table 2 gives an overview of the measured speedups in terms of median runtimes without optimistic estimate pruning divided by median runtimes with optimistic estimate pruning. Absolute runtimes and experiments with different pattern lengths are reported in Section B.5.
The results show substantial runtime reductions. The most prominent improvements were observed for ARL with speedups up to $538.9$ on the Credit dataset, reducing runtime from $\sim 1.25$ hours to less than 10 seconds. For PR AUC, the highest speedup of $22.6$ was also on Credit with runtime reductions from $\sim 1.25$ hours to $\sim 4$ minutes. Finally, for ROC AUC, a speedup of $101.0$ on Census was the highest with an improvement from $\sim 20$ hours to $\sim 12$ minutes. Overall, larger datasets (both in terms of number of instances and features) benefit the most from optimistic estimates across all performance measures. For the smallest datasets (with generally small runtimes) like Approval, where the search split contains less than 250 instances, the speedups substantially decrease across all measures, leading to an overall slowdown in some cases. With regard to parameterization, a very low emphasis on size and class imbalance compared to the performance measure ($\alpha=\beta=0$) as well as a high emphasis ($\alpha=\beta=1$) seems to produce the highest speedups. Further investigations are required, however, an explanation for this may be that the search can more strongly focus on one or the other optimistic estimate factor
@@ 4.2 Optimistic Estimates Substantially Speed Up Runtimes
Table 2: Speedups of median runtimes using optimistic estimates. Speedups are given in terms of median runtimes without optimistic estimate pruning divided by median runtimes with optimistic estimate pruning. Substantial runtime reductions were achieved on all but the smallest dataset. (Cont. in Supp. Table 12, containing results for more datasets)
@@
ARL 0    18.4 47.4 9.7 538.9 1.7
0.1    18.6 39.4 7.7 385.6 0.8
0.3    19.9 40.4 7.6 383.1 0.7
1    23.3 44.7 7.6 388.4 0.6
PR AUC 0    2.9 2.3 1.7 2.5 1.2
0.1    1.9 1.2 0.7 1.1 0.4
0.3    2.7 2.1 0.8 2.9 0.4
1    11.4 19.9 2.2 22.6 0.4
ROC AUC 0    101.0 0.9 2.1 83.8 0.8
0.1    1.0 0.8 0.8 1.0 0.5
0.3    1.3 1.0 0.9 1.1 0.5
1    6.4 8.4 2.1 3.9 0.5
4.3 Discovering Injected Subgroups
To investigate the validity of SubROC and its ability to discover expected subgroups, we injected an artificially underperforming subgroup into a dataset based on the predictions of an xgboost classifier on the Adult dataset. The injected subgroup was selected randomly from all subgroups with pattern length $\leq 3$ in a cover size range from $0.4\%$ to $0.6\%$ of the search split. It covers 54 instances, has an NCR of 0.2, and an ARL of 1.07, PR AUC of 0.97, and ROC AUC of 0.90 before and 9.93, 0.62, and 0.1, respectively, after injection. For the injection, we multiplied the predicted scores of instances in the subgroup cover by $-1$. Thereby, the injected subgroup forms a known weakness in the predictions, and we can evaluate the subgroup search by whether this subgroup is included in the result set. The subgroup discovery was performed up to depth 4 under different configurations (baseline without the SubROC framework and three different weightings of cover size and class balance in the SubROC framework). Result sets were top-10 filtered to reflect practically interpretable result set sizes.
Figure 2 shows the IoU as an overlap measure between found subgroups and the injected subgroup for the ROC AUC. In contrast to the baseline, the SubROC framework successfully discovers the injected subgroup and reports it at the highest ranks in the result set. Note that multiple patterns can describe the same set of instances, so the injected errors can be found several times.
While ROC AUC robustly identifies the subgroup, similar experiments on ARL or the PR AUC did not reproduce these findings. The result sets contain only subgroups with IoU s with the injected subgroup below $0.1$. See Section B.4 for detailed results. Differences in terms of cover size and class balance skew (see Section 4.4) between the performance measures have a large influence on the results of these experiments. The injected subgroup contains more negatives than positives, which interacts with the class balance skews of ARL and PR AUC. Furthermore, the injected subgroup is small, making it inherently hard to capture with ARL, which has an embedded size skew. This highlights the importance of the size and class balance terms introduced in SubROC to counter these skews, but also shows that they do not completely remove the effects. Future work, for example, employing a normalized PR AUC [boyd_unachievable_2012] may be promising.

### v2

@@ 3.1.3 Generalization Awareness
It is a common issue in subgroup discovery that multiple subgroups in the result set have a high pairwise overlap in terms of their cover [20]. To reduce redundancy along the generalization-specialization hierarchy, adaptations to scoring functions have been proposed under the term generalization awareness. In this approach, a variant [20] of previous approaches [21, 22, 23] extends a scoring function $\varphi^{\prime}$ by reducing the score $\varphi^{\prime}(p)$ of a subgroup with pattern $p$ by subtracting the maximum score across all its generalizations $h$ (making a specialization viable only if its score is substantially higher):
@@ 3.2 Optimistic Estimates
For runtime optimization, we derive tight optimistic estimates for the scoring functions $\varphi^{rasl}$, $\varphi^{rROCAUC}$ and $\varphi^{rPRAUC}$ as well as, separately, for the cover size and class balance factors. Proofs are given in Appendix˜A.
@@ 3.2.1 Average Ranking Loss
To date, no (tight) optimistic estimate was derived for the ARL-based scoring function $\varphi^{rasl}$ [8]. $ARL(p)$ is defined as the average over the ranking losses $PEN(c,p)$ of all covered positive instances $c\in\mathcal{P}_{p}$. The penalty $PEN(c,p)$ for each positive instance $c$ counts the negative instances ranked higher than $c$. Thus, removing positive instances with non-maximal $PEN(c,p)$ does not alter the $PEN$ of the remaining instances and increases the average by eliminating small values. The $ARL(p)$ is therefore maximized when only a single positive $c^{\prime}$ instance with maximal $PEN(c^{\prime},p)$ is left.
@@ 3.2.2 Area Under the Receiver-Operating Curve
Next, we introduce a lower bound $b$ for ROC AUC that we will use to derive an optimistic estimate of $\varphi^{rROCAUC}$. We distinguish three cases: (a) If perfect separation of ground truth labels is possible with a threshold on the soft predictions, the same holds for every subset of the cover. Thus, the ROC AUC is always 1. (b) If the input contains some instances that share the same prediction value (ties) but have different labels (positive and negative), then the ROC AUC on only these tied instances equals $\frac{1}{2}$. Any subset of the given instances containing both classes and at least one non-tied instance allows for the correct separation of some positives and negatives by at least one threshold. Thus, the ROC AUC is greater than $\frac{1}{2}$ for such subsets. (c) If the input contains at least one incorrectly ranked pair of a positive and a negative instance. The subset containing exactly one such pair has a ROC AUC of 0, which is the lowest possible ROC AUC.
@@ 3.2.3 Area Under the Precision-Recall Curve
Lemma˜4 introduces a lower bound for PR AUC which is used in Theorem˜5 to derive a tight optimistic estimate for $\varphi^{rPRAUC}$. For Lemma˜4, a subset $I_{worst}$ of the original set of instances $I$ is constructed to maximize classification errors while minimizing the positive class ratio. This makes it impossible for a threshold on any other subset to define a point in PR space below the PR curve of $I_{worst}$ (with some exceptions that have no impact on the AUC). Proofs and an illustration are included in Section˜A.3.
@@
3.2.4 Cover Size and Class Balance Weighted Scoring
The following Lemma˜6 gives a tight upper bound for the weighting used in cover size and class balance weighted scoring functions when $\alpha\leq\beta$. This result translates to optimistic estimates for $\varphi_{\alpha,\beta}^{rasl}(\cdot)$, $\varphi_{\alpha,\beta}^{rROCAUC}(\cdot)$ and $\varphi_{\alpha,\beta}^{rPRAUC}(\cdot)$. It therefore enables the use of the optimistic estimates presented in theorems 1, 3 and 5 when cover size and class balance weighting is enabled.
@@ 3.3 Significance Filtering
Subgroup discovery inspects an exponential search space of patterns and consequently requires correction for multiple hypothesis comparison [6]. To reduce the number of compared subgroups, we use a holdout procedure [24]: We split the dataset into two parts, the search dataset to search subgroups and the validation dataset to test statistical significance. The tested statistic is the raw score of a subgroup (written $\varphi_{0,0}$) excluding cover size and class imbalance. The null hypothesis is that $\varphi_{0,0}(p)$ for a pattern $p$ is drawn from the sample distribution of $\varphi_{0,0}$ on random instances with the same size and class balance as the cover $sg(p)$ of $p$. This allows a test on performance differences, excluding the influence of cover size and class balance on the used performance measure.
We compute the empirical p-value of a subgroup $p$ by a randomization test (cf. [25]), where $1000$ random subsets of the validation dataset are drawn without replacement. To prevent significance due to size and class balance, we keep these parameters constant during sampling. The empirical p-value is the fraction of random subsets with interestingness $\varphi_{0,0}$ greater than or equal to $\varphi_{0,0}(p)$ divided by the total number of random subsets.
Multiple testing correction is done using Benjamini-Yekutieli (BY) [26] (controlling the false discovery rate (FDR)). BY has the necessary properties to be applied under any dependency between tested hypotheses. The correction approach can be customized by the user, e.g., with Bonferroni controlling familywise error rate (FWER). The default p-value threshold we use is $p\leq\alpha$. Here $\alpha=0.05$.
This enables significance filtering, focusing the search on significant subgroups: To find top-$k$ significant subgroups (e.g., $k=5$), search for the top-$k^{\prime}$ subgroups with $k^{\prime}\gg k$ (e.g., $k^{\prime}=100$), calculate p-values, and only include a subgroup in the final top-5 if their corrected p-value passes the significance threshold $\alpha$.
@@ 4 Experiments
Datasets are from OpenML [27] and the UCI Machine Learning Repository [28] (statistics in Table˜4). Each dataset was partitioned randomly into three equal parts: (i) a training set for fitting the prediction model; (ii) a search set where the trained model is applied to yield predictions $\hat{Y}$ on unseen data and the subgroup discovery is performed; (iii) a validation set for significance testing of the found subgroups (see Section˜3.3).
4.1 Quality of SubROC Across Multiple Datasets
Table 1: Overview of top-5 results from a basic SubROC setting (b) and the full SubROC framework (f) for a range of datasets and performance measures. Full SubROC managed to find diverse and exceptional subgroups, which are often much larger and balanced than the baseline. Numbers are bold for the better result when comparing b and f. (Cont. in Table˜5, containing results for more datasets)
@@ 4.1 Quality of SubROC Across Multiple Datasets
Filtered-5 / Significant / Top-100
Mean Exception- ality
@@
Dataset Metric     b f     b f     b f     b f     b f
Census n=199523 NCR=0.06 ROC     25 6753    1 / 5 5 / 99 / 100    0.95 0.05    0.04 0.16    0.10 0.31
PR     23 864    0 / 5 5 / 99 / 100    0.92 0.14    0.93 0.45    0.37 0.45
ARL     8000 11945    5 / 5 5 / 27 / 27    47 31    0.15 0.11    0.99 0.50
Adult n=48842 NCR=0.24 ROC     27 6492    0 / 5 5 / 77 / 100    0.92 0.06    0.04 0.37    0.48 0.61
PR     27 2803    2 / 5 5 / 70 / 100    0.94 0.08    0.96 0.46    0.37 0.19
ARL     4614 6930    5 / 5 5 / 5 / 5    148 55    0.42 0.34    0.85 0.61
Bank n=45211 NCR=0.88 ROC     22 311    0 / 5 5 / 22 / 100    0.83 0.16    0.95 0.48    0.10 0.12
PR     157 72    0 / 5 2 / 2 / 100    0.68 0.10    0.99 0.48    0.80 0.51
ARL     - -    0 / 0 0 / 0 / 0    - -    - -    - -
Credit n=30000 NCR=0.78 ROC     24 1506    0 / 5 5 / 38 / 100    0.75 0.12    0.96 0.61    0.01 0.09
PR     75 1895    0 / 5 5 / 17 / 100    0.51 0.29    0.99 0.85    0.18 0.34
ARL     - -    0 / 0 0 / 0 / 0    - -    - -    - -
Approval n=690 NCR=0.55 ROC     29 -    0 / 5 0 / 0 / 41    0.64 -    0.96 -    0.69 -
PR     41 80    2 / 5 1 / 1 / 25    0.89 0.85    0.98 0.95    0.75 -
ARL     60 87    3 / 5 3 / 3 / 3    26 10    0.96 0.79    0.68 0.69
Table 2: Example filtered-5 result set from the full SubROC framework using ROC AUC on Adult data with xgboost predictions.
Pattern ROC AUC Cover NCR
$\text{marital-status}=\text{Married-civ-spouse}$
0.8508
5280
0.4496
$\text{relationship}=\text{Husband}$
0.8485
4638
0.4495
$\text{capital-gain}\in[0.0,114.0)\>\wedge$
$\text{marital-status}=\text{Married-civ-spouse}$
0.8053
4628
0.4114
$\text{sex}=\text{Male}$
0.9053
7581
0.3109
$\text{capital-gain}\in[0.0,114.0)$
0.9015
10331
0.2126
$\emptyset$
0.9241
11305
0.2482
To illustrate the effectiveness of the full SubROC framework, we compare a basic SubROC setting (b) and the full SubROC framework (f) for finding the top-5 underperforming subgroups (Table˜1). Both cases train xgboost [29] as the binary classifier on (i), chosen because it is a widely adopted, strong baseline method. Subgroups are then exhaustively [^2] Results are the same for any non-heuristic (exhaustive) search algorithm. searched on (ii) with a minimum subgroup size of 20, and a maximum pattern length of 4. The basic setting (b) solely uses the bare scoring functions $\varphi^{rasl}$, $\varphi^{rROCAUC}$, and $\varphi^{rPRAUC}$ and returns the top-5 subgroups. Their significance is tested on (iii) to underpin their tendency for random artifacts (e.g. very small subgroups, high class imbalance), see column Significant / Top-5. The full framework (f) is configured with size and class imbalance weighting $\alpha=\beta=1$, generalization awareness, and statistical significance filtering ($k^{\prime}=100$). To illustrate the effectiveness of significance filtering, we report how many subgroups were found for a top-100 search on (ii), how many of these passed the statistical test on (iii), and how many finally remain after significance filtering (see column Filtered-5 / Significant / Top-100). Statistics were calculated for the final top-5 results on (ii) in both cases. Note that exceptionality should not be compared between different performance measures because of their differing scales and responses to errors.
For b, cover sizes tend to be close to the minimum cover size limit, and class ratios tend to be close to 0 or 1. Despite their high difference in performance compared to the overall dataset (exceptionality), such subgroups do not necessarily point towards structural issues of the classifier but may be caused by artifacts (cf. Section˜4.4). This is also apparent in the low number of subgroups passing statistical significance in the top-5 results of b. Similar patterns are observable in the top-100 results (see Table˜6). While yielding less prominent exceptionality, we observe that the SubROC framework leads to larger and more balanced results. Additionally, returned subgroups are substantially more likely to pass the significance filtering step.
Mean pairwise Intersection over Union (IoU) of the discovered subgroups was computed to assess redundancy. Values for SubROC remain comparable to the baseline, with some significant improvements, indicating that our framework does not sacrifice diversity.
The significance filtering step of SubROC saves resources for practitioners by discarding a considerable amount of initially found, but unreliable subgroups. This effect is particularly prominent for small datasets with less than $2,500$ instances in the search set, where the significance filtering discarded almost all subgroups. It is less prominent for ARL based scoring functions, which generally produce smaller result sets with a greater share of significant subgroups.
Table˜2 shows one of the result sets from this experiment. For more full result sets on the Adult dataset, see Section˜B.3.
4.2 Runtime Speed Ups with Optimistic Estimates
@@
Table 3: Speedups of median runtimes using optimistic estimates, given in terms of median runtimes without optimistic estimate pruning divided by median runtimes with optimistic estimate pruning. (Cont. in Table˜13, containing results for more datasets)
@@
ARL 0    9.1 11.6 482.0 10.2 49.3
0.1    9.3 11.1 333.9 9.3 24.7
0.3    10.4 13.0 337.6 9.5 24.6
1    15.0 27.5 333.3 9.3 25.4
PR AUC 0    2.0 1.5 460.7 3.3 43.8
0.1    2.0 1.2 337.2 1.7 25.4
0.3    2.0 1.2 338.0 1.6 24.6
1    3.8 5.3 337.1 2.2 24.9
ROC AUC 0    1.5 1.1 401.7 1.7 41.7
0.1    1.5 1.0 332.8 1.6 17.0
0.3    1.4 1.0 303.2 1.6 19.5
1    2.8 4.4 307.4 1.8 24.2
To investigate the efficiency improvements enabled by SubROC, we measure runtimes of subgroup discovery with and without pruning based on our optimistic estimates, carrying out 10 repetitions for each dataset (3 repetitions for Census due to long runtimes). The searches for the top-5 patterns run on predictions from an xgboost classifier, using the best-first-search enumeration strategy, with a maximum pattern length of 4. Different size and class balance weightings ($\alpha,\beta$) are compared. Table˜3 gives an overview of the measured speedups. Absolute runtimes are given in Section˜B.5.
The results show substantial runtime reductions. Particularly large speedups were achieved on the Bank dataset. For ARL, pruning cut the runtime from $\sim 2.13$ min to <0.3<0.3 s ($\sim 482\times$ speedup). For PR AUC, the runtime dropped from $\sim 2.42$ min to $0.3$ s ($\sim 461\times$ speedup). For ROC AUC, pruning reduced the runtime from $\sim 2.67$ min to $0.4$ s ($\sim 402\times$ speedup). With regard to parameterization, both ignoring size and class balance altogether ($\alpha=\beta=0$) and weighting them equally to the performance measure ($\alpha=\beta=1$) appear to produce the highest speedups. Further investigations are required, however, an explanation for this may be that the search can more strongly focus on one or the other optimistic estimate factor.
4.3 Discovering an Injected Subgroup
(a) baseline
(b) $\alpha=0,\beta=0$
(c) $\alpha=0.3,\beta=0$
(d) $\alpha=0,\beta=0.3$